OpenStudy (anonymous):

At the start of the basketball game, the referee tosses a ball for a jump ball. The equation that models the height of the ball is h(t) = -16t2 + 24t + 5 During what time interval is the height of the ball above 9 feet? (Round your answer to the nearest tenth of a second.) a) .5 seconds to 2.1 seconds b) .4 seconds to 1.9 seconds c) .6 seconds to 1.8 seconds d) .2 seconds to 1.3 seconds

OpenStudy (anonymous):

4. h(t) = -16t^2 + 30t + 8 10=-16t^2 + 30t +8 0= -16t^2 + 30t -2 Zeros at t= 0.069 and t=1.81 using the quadratic equation. c) Between 0.1 and 1.8 seconds